OpenStudy (anonymous):

If h(x) = 5 for the function h(x) = 2x + 1, what is the value of x?

OpenStudy (anonymous):

\(5=h(x) = 2x+1\), so solve \(5=2x+1\).

jhonyy9 (jhonyy9):

2x+1=5 2x=4 x=2

OpenStudy (anonymous):

thanks you wingspan lover